the sum of the ages of a father and his son is 41years If the product of their ages is 210 find their ages I want that answer now urgently
Answer by rothauserc(4718)

let x be the father's age and y be the son's age, then
:
1) x + y = 41
:
2) xy = 210
:
solve equation 1 for x and substitute for x in equation 2
:
x = 41 - y
:
(41-y)y = 210
:
41y -y^2 = 210
:
y^2 -41y +210 = 0
:
(y-35) * (y-6) = 0
:
y = 35 or 6
:
there are two possible answers
:
1) father is 35 and son is 6
:
2) father is 6 and son is 35
:
we reject answer 2
:
****************************************
father is 35 and son is 6
****************************************
: